A menu at a restaurant offers 9 different appetizers. How many different ways can a group choose 4 appetizers?
Answer by jim_thompson5910(35256)

4 Slots: A,B,C,D

Each slot represents an appetizer selected

Number of choices for slot A = 9

Number of choices for slot B = 8 ... this is assuming you cannot have repeat appetizers

Number of choices for slot C = 7

Number of choices for slot D = 6

--------------------------------------------------------------

Multiply them out: 9*8*7*6 = 3,024

If order mattered, then you'd have 3,024 different permutations

Order does matter though. So divide by 4! = 4*3*2*1 = 24 to get 3,024/24 = 126

There are 126 different combinations (where order does not matter)

--------------------------------------------------------------

Question: A menu at a restaurant offers 9 different appetizers. How many different ways can a group choose 4 appetizers?

Answer: 126


------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------